# Quantum Critical Behavior in Ce-Au-Al Quasicrystal Approximants

### v1

@@ IV Conclusions
We have studied the effect of magnetic field on the low-temperature magnetic susceptibility and specific heat capacity for the 1/1 and 2/1 ACs of the Ce-Au-Al system. The systems display diverging magnetic susceptibility at zero fields and an energy-scale competition between temperature and magnetic field that set the cut-off of critical fluctuations. This strongly suggests that these ACs show a resemblance with quantum critical systems Löhneysen et al. (2007). Specific heat does not find any magnetic ordering, but also does not show any clear signatures of critical behavior. Further investigations are needed to rule out magnetic order below 0.2 K. The alternative ground state for both AC systems would be a quantum paramagnetic spin-liquid state Gao et al. (2019), unless the systems are located right at the quantum critical point.

### v2

@@ IV Conclusions
We have studied the effect of magnetic field on the low-temperature magnetic susceptibility and specific heat capacity for the 1/1 and 2/1 ACs of the Ce-Au-Al system. The systems display diverging magnetic susceptibility at zero fields. Even though the measured ac-susceptibility in magnetic field closely resembles the behavior of a quantum critical system, the major characteristics can be explained by the splitting of the ground state Kramers doublet in magnetic field. The reason for this resemblance is that the Kramers doublet acts as a precursor for the heavy fermion quantum criticality. Specific heat does not find any magnetic ordering, but also does not show any clear signatures of critical behavior. Further investigations are needed to rule out magnetic order below 0.2 K. We anticipate that stronger interaction of conduction electrons with the Kramers doublet will lead to modification of both the exponent $\beta$ and the $g$-factor $g_{\mathrm{K}}$ of the doublet such that it develops a strong temperature dependence. The corresponding redistribution of entropy will be an indication of critical fluctuations and a quantum critical ground state. By substituting Ce with Yb, stronger interactions are induced, and a quantum critical state is achieved, both in the ACs and QCs [12, 11]. Partial substitution of this sort is anticipated to illustrate how the current precursor state becomes quantum critical.